An actor invested some money at 6​% simple​ interest, and ​$44,000 more than four times the amount at 12 %. The total annual interest earned from the investment was ​$50,640. How much did he invest at each​ amount?
Answer by Boreal(15235)

x at 6%
4x+44000 at 12%
.06x+(0.12)(4x+44000)=50640
.06x+.48x+5280=50640
.54x=45360
x=$84000 at 6%, and $5040 interest
4x+44000=$380000 at 12% and $45,600 interest